# LionHeart: A Layer-based Mapping Framework for Heterogeneous Systems with Analog In-Memory Computing Tiles

### v1

@@ IV-B Training and Evaluation Strategies
IV-B1 Initial hyper-parameter exploration:
@@
Our framework assumes that the input DNN is pre-trained in floating-point precision. The selected models differ in depth, size, and complexity, and hence require different hyperparameters to achieve optimal training performance. We empirically determined two sets of hyper-parameters, pertaining to the analog and digital domain, respectively. Then, these were applied to all networks. For analog training, a learning rate of 0.024 and momentum of 0.775 was used. In terms of digital training, we employed a learning rate of 0.057 and a momentum of 0.867. A convergence window of 5 epochs was used for retraining. Further details on the adopted experimental setups in the two domains are provided in the following.
IV-B2 Digital training:
@@
Networks were trained using floating-point precision until the training loss did not decrease for five consecutive epochs. The batch size was set to 256. Stochastic Gradient Descent (SGD) with momentum was used, with cross-entropy loss. A cosine-annealing learning rate scheduler was used.
IV-B3 Analog HWA re-training:
@@
To perform analog HWA re-training, the weights of the trained digital networks were employed as a starting point. The IBM AIHWKIT was used to inject noise during forward propagations. We refer the reader to [6] for a comprehensive tutorial on HWA training using IBM AIHWKIT. The InferenceRPUConfig() preset was used with the PCMLikeNoiseModel phenomenological inference model. The following additional modifications were made to the simulation configuration: (i) biases are assumed to be digital (i.e., they are not encoded on the last column of AIMC tiles). (ii) Channel- (i.e., column-) wise weight scaling, which has a negligible performance impact, is performed, where weights are mapped to a per-channel conductance range during learning after each mini-batch. (iii) The size of each AIMC tile is assumed to be 256x256. (iv) Layers that span across multiple AIMC tiles are assumed to be split evenly. (v) During training, multiplicative Gaussian noise is applied to unit weight values, with a standard deviation of $0.08$, extracted from hardware measurements provided by the authors of [4].
IV-B4 Accuracy evaluation:
@@
As AIMC hardware is inherently stochastic, multiple evaluation instances are required to determine representative performance metrics. Consequently, accuracy was measured 20 times and averaged across experiments to obtain the values reported in Section IV-D. We considered five user-defined accuracy thresholds of [0.5%, 1.0%, 3.0%, 5.0%, 10.0%]. Moreover, analog NVM devices such as PCM are susceptible to temporal conductance drift. Since the behavior of AIMC hardware can evolve over time, we explore in Section IV-F accuracy metrics for AIMC accelerators programmed considering drift levels predicted at [one second (i.e., $t=1s$), one minute (i.e., $t=60s$), one hour (i.e., $t=3600s$), one day (i.e., $t=86400s$)]. These are then evaluated at different points in time, to assess the effect of a mismatch between predicted and actual drift.
IV-B5 Full System Evaluation:
@@
The target simulated system for this work is a high-performance edge processor with an in-order ARM CPU core running at a frequency of 2.3GHz, with a 64KB each L1 data and instruction caches, 1MB of L2 cache, and 8GB of 2400MHz DDR4 RAM. The software stack for the target DNNs includes the DNNs accelerated via the Eigen C++ library, running in user-space atop the Linux 5.4 kernel, and an Ubuntu 16.04 LTS disk image. Each implementation resulting from a LionHeart mapping is run for 10 inferences to reduce simulation noise. We note that measured deviations among different runs of the same mappings never exceeded 4% of the run-time.
@@ IV-C Baselines
LionHeart (in all plots, referred to as LH) is evaluated by comparing the achieved accuracy and performance with respect to three baselines: (i) Fully-digital. The floating-point DNN model input of our methodology, where none of its layers are accelerated using AIMC. It achieves the highest accuracy, but presents the longest run-time. (ii) Fully-analog. A DNN implementation with all FC and CONV layers executed using AIMC. This baseline represents the opposite of the fully-digital: it maximizes performance, while potentially resulting in very large accuracy degradation. (iii) First-Last Mapping Strategy (FLMS). The SOTA hybrid mapping strategy [11]. In these mappings, only the first and the last layers are implemented in digital PUs, while all others are AIMC-accelerated.
@@ IV-D Accuracy Evaluation
For five different drop thresholds and baseline configurations, we determined the ratio between MACs performed using AIMC and the total number of MACs (Fig. 3). As can be observed, for all networks, the All Digital configuration has the highest test set accuracy and the All Analog configuration has the lowest test set accuracy. The behavior of the FLMS configuration, however, is unpredictable. The proposed LionHeart configurations exhibit a gradual transition from the largest to the smallest test set accuracy and a corresponding decrease in the digital/analog MAC ratio between all configurations. It hence enables the exploration of the performance/accuracy space exposed by AIMC acceleration, while controlling accuracy to a user-defined constraint.
For 62% of configurations, the last FC layer is retained in analog. However, for only 18%, the first CONV layer is retained. This means that the first layer is usually the most sensitive. Intermediary layers are observed to be more sensitive than the last layer. For intermediary layers, no strong trend or indication is observed as to whether a layer will remain analog, other than that all FC ones are more likely to be retained as digital layers. Hence, it is likely, for CONV layers, that whether or not a layer will be retained in analog is largely dependent on the trained weights and not on other layers’ characteristics. Our analyses demonstrate that properly mapping layers to digital or analog resources is non-trivial, showcasing the benefit of our framework in comparison with more naïve strategies such as FLMS.
IV-E Performance Evaluation
TABLE II: Architectural parameters for full system simulation.
@@ IV-E Performance Evaluation
Figure 4: (a) Speedup and (b) energy efficiency gain resulting from different application mapping strategies, when performing an inference on benchmark DNNs. Data is normalized with respect to a fully digital execution with no AIMC acceleration.
@@
In LionHeart, the workload associated with DL models and their layers is assumed to be architecture-agnostic, i.e., the ratio of analog MAC to total MAC operations is used as a proxy for speedup and energy efficiency. To quantify how strongly the MAC ratio correlates to these metrics, we evaluate them for a simulated system featuring a single-core in-order CPU clocked at 2.3GHz and a two-level cache hierarchy. The architecture embeds tightly-coupled AIMCs, governed by dedicated extensions to the ARM instruction set supported by the CPU. It is modeled using the ALPINE framework [15], allowing the full system simulation of AIMC-accelerated systems. Table II lists the adopted architectural parameters, which mirror the ones considered in [15] for the single-core, higher-performance scenario.
Due to the L1 and L2 cache sizes (rendering AIMC-acceleration of smaller networks ineffective), results were gathered on the three largest DNNs among the ones presented in Section IV-D, considering the more challenging CIFAR-100 dataset. We simulated the execution of an entire inference, including computations not pertaining to FC nor CONV layers. We investigated heterogeneous mappings derived by LionHeart, with varying maximum accuracy degradation thresholds. For comparison, we also conducted experiments where all FC and CONV layers are completely executed on AIMCs, and mappings where, similar to [11], all layers except the first and last are hosted on analog crossbars (FLMS).
@@
The complete simulation outcomes of the system, presented in Fig. 4, highlight that the achievable speed-ups are highly dependent on the DNN structure. In particular, even when all layers are computed in the analog domain, which is an upper bound for speedup and efficiency in our setting, only a speedup of 24% can be reached in the MobileNet benchmark. Such (comparatively) low run-time performance improvement is caused by the lower degree of data reuse of the depthwise separable convolutions employed in this network, which shift the balance between computation and memory access requirements toward the latter, limiting the benefit of accelerating MACs. Nonetheless, even in this case, LionHeart, with a limited maximum accuracy degradation of 5%, still captures almost all of the available speedup and energy efficiency gains.
@@
AlexNet and VGG16, which do not employ separable convolutions, present instead more potential speedup (660% in the full-analog case for VGG16, 550% for AlexNet), due to the higher locality of the employed computations (which leads to less cache misses and memory-induced stalls). Even in those cases, LionHeart can harness a large part of the potential speedup if enough accuracy leeway is allowed. For a 5% accuracy degradation, a 425% speedup was measured in AlexNet and a 285% one in VGG16. AlexNet proved to be the most robust of the investigated networks. In fact, for a 10% accuracy degradation threshold, its performance exceeded that of the FLMS baseline, as all layers can be executed in AIMCs, further showcasing the benefit of adopting a flexible accuracy-driven approach. Finally, it can be noticed that system-level energy gains closely follow speedups. Indeed, AIMCs tiles are not a major contributor to overall system power (as opposed to processors and memories). Instead, they shorten run-time, ultimately lowering energy requirements. This is achieved both by allowing the execution of MVMs in constant time and by encoding DNN weights in the crossbar structure itself, hence easing the pressure on the memory hierarchy.
IV-F Temporal Drift and Evaluation Time Analysis

### v3

@@ IV-B Training and Evaluation Strategies
IV-B1 Initial hyper-parameter exploration
@@
Our framework assumes that the input DNN is pre-trained in floating-point precision. The selected models differ in depth, size, and complexity, and hence require different hyperparameters to achieve optimal training performance. For all networks listed in Table I, we empirically determined two sets of hyper-parameters, pertaining to the analog and digital domain, respectively. Then, these were applied to the networks. These are listed in Table II. Further details on the adopted experimental setups in the two domains are provided in the following.
IV-B2 Digital training
@@
For all Convolutional Neural Networks, networks were trained using floating-point precision until the training loss did not decrease for five consecutive epochs. For MobileBERT, a fixed number (3) of training epochs was used. All networks were trained with cross-entropy loss in conjunction with Stochastic Gradient Descent (SGD) with momentum.
IV-B3 Analog HWA re-training
To perform analog HWA re-training, the weights of the trained digital networks were employed as a starting point. The IBM AIHWKIT was used to inject noise during forward propagations. We refer the reader to [5] for a comprehensive tutorial on HWA training using IBM AIHWKIT. To configure the behavior of AIMC tiles and HWA training, the IBM AIHWKIT utilizes different configurations which are defined using a standardized data class data structure. A number of pre-defined configurations are provided. We used the provided InferenceRPUConfig() configuration and PCMLikeNoiseModel phenomenological inference model.
@@ IV-B3 Analog HWA re-training
The following additional modifications were made to the simulation configuration: (i) biases are assumed to be digital (i.e., they are not encoded on the last column of AIMC tiles). (ii) Channel- (i.e., column-) wise weight scaling, which has a negligible performance impact, is performed, where weights are mapped to a per-channel conductance range during learning after each mini-batch. (iii) The size of each AIMC tile is assumed to be 256x256. (iv) Layers that span across multiple AIMC tiles are assumed to be split evenly. (v) During training, multiplicative Gaussian noise is applied to unit weight values, with a standard deviation of $0.08$, extracted from hardware measurements provided by the authors of [12].
IV-B4 Accuracy evaluation
@@
As AIMC hardware is inherently stochastic, multiple evaluation instances are required to determine representative performance metrics. Consequently, accuracy was measured 20 times and averaged across experiments to obtain the values reported in Section IV-D. We considered five user-defined accuracy [^4] For SQuAD, instead of considering the accuracy, the F1 score is considered as a percentage value. thresholds of [0.5%, 1.0%, 3.0%, 5.0%, 10.0%]. Moreover, analog NVM devices such as PCM are susceptible to temporal conductance drift. Since the behavior of AIMC hardware can evolve over time, we explore in Section VI accuracy metrics for AIMC accelerators programmed considering drift levels predicted at [one second (i.e., $t=1s$), one minute (i.e., $t=60s$), one hour (i.e., $t=3600s$), one day (i.e., $t=86400s$)]. These are then evaluated at different points in time, to assess the effect of a mismatch between predicted and actual drift.
IV-B5 Full System Evaluation
@@
The target simulated system for this work is a high-performance edge processor with an in-order ARM CPU core running at a frequency of 2.3GHz, with a 64KB each L1 data and instruction caches, 1MB of L2 cache, and 8GB of 2400MHz DDR4 RAM, as detailed in Table IV. The software stack for the target DNNs includes the DNNs accelerated via the Eigen C++ library, running in user-space atop the Linux 5.4 kernel, and an Ubuntu 16.04 LTS disk image. Each implementation resulting from a LionHeart mapping is run for 10 inferences to reduce simulation noise. Measured execution time deviations among different runs of the same mappings never exceeded 4% of the run-time.
@@ IV-C Baselines
LionHeart (in all graphs, referred to as LH) is evaluated by comparing the accuracy and performance achieved with respect to two baselines: (i) Fully digital. The floating-point DNN model input of our methodology, where none of its layers is accelerated using AIMC. It achieves the highest accuracy, but presents the longest run-time. (ii) Fully-analog. A DNN implementation executing all FC and CONV layers with static weights using AIMC. This baseline represents the opposite of the fully-digital: it maximizes performance, while potentially resulting in very large accuracy degradation. In Section V, we provide a further comparative evaluation with state of the art strategies, namely Ueyoshi et al. [7] (i.e., DIANA) and Harmonica [8].
@@ IV-D Accuracy Evaluation
For five different drop thresholds and baseline configurations, we determined the ratio between MACs performed using AIMC and the total number of MACs for CIFAR-10, CIFAR-100 and SQuAD (see Fig. 5). As expected, for all networks the All Digital configuration has the highest test set accuracy and the All Analog configuration has the lowest test set accuracy. The proposed LionHeart configurations exhibit a gradual transition from the accuracy of the largest to the smallest test set and a corresponding decrease in the digital/analog MAC ratio between all configurations. Thus, it enables exploration of the performance/accuracy space exposed by AIMC acceleration. Critically, the design space is navigated in a controlled way, that is, abiding to a user-defined accuracy degradation constraint. In this way, configurations exhibiting both high performance (high ratio of AIMC acceleration) and high accuracy can be derived. As an example, almost 60% of the computation can be AIMC-accelerated in the ResNet-8 network for an accuracy threshold of 5%, while an all-analog alternative would have a decrease in accuracy of 20%.
For a small number of networks and datasets, it is observed that the All Analog configuration has a higher test set accuracy than some of the LionHeart configurations; particularly those with larger drop thresholds, e.g., 10%. Moreover, some LionHeart configurations with larger drop thresholds have a higher accuracy than others with smaller drop thresholds. To this end, we make the following observations. First, heuristic-based methods are inherently stochastic in outcome. While we investigate the variability of LionHeart and related work in Section V, in this section, we first investigate the one-shot behavior of LionHeart in earnest. Second, LionHeart enforces a global drop threshold, meaning that if a network fails to converge when a layer is retrained using HWA training, a large accuracy drop can be incurred (as large as this threshold). This can result in a sub-optimal outcome if one of the first layers incurs a large accuracy drop, as all subsequent layers are subjected to the remaining drop threshold.
Lastly, given a fixed accuracy threshold, LionHeart does not penalize configurations for being close to this threshold. This means that while in practice it would not make sense to specify, if an accuracy drop of Xwas indeed specified, and the accuracy drop of the fully-analog configuration was X-N%, the resulting mapping would not be penalized for being between X-Nand X%.
TABLE III: Ratio of configurations where a layer is AIMC-accelerated, per layer type and topological location in the MobileBERT encoder.
Encoder Block Location First Middle (Avg.) Last
Encoder Layer Average Retention in Analog (%)
Bottleneck 45 52 46
Attention 78 81 61
FFN0 80 68 56
FFN1 75 74 62
FFN2 68 70 66
Intermediate 48 75 48
Output 45 52 40
We observe that no obvious pattern emerges regarding the topological order of layers in our explorations, other than that CONV layers are more likely to be AIMC-accelerated compared to FC ones. For the same layer type, whether or not a layer will be executed in the analog domain depends instead on the trained weights, and not on the layers and the network topology, thus mandating exploration heuristics such as LionHeart. A similar conclusion can be reached for the MobileBERT case. Again, the choice of accelerated layers does not follow a clear topological pattern, as we report in Table III, where the conversion percentage is reported for the first and last, in addition to the average of all middle encoder blocks. The first attention layers are more amenable to be executed in AIMCs, but this is not the case for Bottleneck layers.
From our analysis, for both CNNs and transformer-based networks, we observe that the first layer is usually one of the most sensitive. The intermediate layers are generally seen to be more sensitive than the last layer. Our analyses demonstrate that, especially for larger networks with a more complex topology, properly mapping layers to digital or analog resources is non-trivial.
TABLE IV: Architectural parameters for full system simulation.
@@
Figure 6: (a) Speedup and (b) energy efficiency gain resulting from different application mapping strategies, when performing an inference on benchmark DNNs. Data is normalized with respect to a fully digital execution with the ARM NEON SIMD co-processor in lieu of AIMC acceleration.
IV-E Performance Evaluation
In LionHeart, the workload associated with DL models and their layers are assumed to be architecture-agnostic, i.e., the ratio of analog MAC to total MAC operations is used as a proxy for speedup and energy efficiency.
This approach allows to derive hybrid mappings in a hardware-agnostic way, while also avoiding impractically lengthy simulations. On the other hand, it does not allow to directly quantify the performance and efficiency of mappings on a target architecture. To inspect this aspect, and also validate the link between MAC ratio and performance/efficiency for one system configuration, we herein execute entire inferences for the three largest CNN benchmarks (AlexNet, VGG16 and MobileNetV2) and measure the ensuing run-times, including computations not pertaining to FC nor CONV layers. We investigate mappings derived by LionHeart, with varying maximum accuracy degradation thresholds.
@@ IV-E Performance Evaluation
As a test vehicle, we adopted a simulated system featuring a single-core in-order CPU clocked at 2.3GHz and a two-level cache hierarchy, modelled within the ALPINE framework [4]. The architecture embeds tightly-coupled AIMCs, governed by dedicated extensions to the ARM instruction set supported by the CPU. Table IV lists the architectural parameters of the test system, which mirror those considered in [4] for the single-core, higher performance scenario.
To align the performance experiments with state-of-the-art low-power edge-domain devices, all DNN implementations are quantized to INT8 precision. The fully digital baseline architecture utilizes the ARM Neon SIMD co-processor in lieu of the tightly-coupled AIMC tiles. Note that due to the effects of memory transactions and latencies, the effective Op/s of both the AIMC-enabled and baseline systems are a fraction of the peak Op/s, at 57% and 35%, respectively.
The CPU in the evaluated system architecture uses customized ISA extensions to access AIMC tiles that are tightly coupled to the CPU cores, without requiring the traversal of the memory hierarchy for data transmission [4]. The extension instructions are implemented using unused opcodes in the ARMv8 architecture. During inference, three instructions are used: CM_QUEUE, which places packed inputs in the input memory of an selected AIMC tile, CM_PROCESS, which operates the AIMC tiles and performs MVMs, and CM_DEQUEUE, which retrieves the outputs from selected AIMC tiles and places them in destination registers. For all speedup and efficiency gain factor evaluations, it is assumed that a single CPU core is used and the required number of tiles to map all analog weights for the evaluated network are tightly coupled to this core.
In Fig. 6, we report the obtained speedups (with respect to fully-digital baselines) and efficiency gain factors. In order to align the performance experiments with state-of-the-art low-power edge-domain devices, all DNN implementations are quantized to int8 precision. Additionally, the fully digital baseline architecture utilizes the ARM Neon SIMD co-processor in lieu of the tightly-coupled AIMC tiles. Note that due to the effects of memory transactions and latencies, the effective Op/s of both the AIMC-enabled and baseline systems are a fraction of the peak Op/s, at 57% and 35% respectively.
@@
These results highlight that the achievable speed-ups are highly dependent on the DNN structure. In particular, even when all layers are computed in the analog domain, which is an upper bound for speedup and efficiency in our setting, only a speedup of 24% can be reached in the MobileNetV2 benchmark. Such (comparatively) low run-time performance improvement is caused by the lower degree of data reuse of the depthwise separable convolutions employed in this network, which shift the balance between computation and memory access requirements toward the latter, limiting the benefit of accelerating MACs. However, even in this case, LionHeart, with a limited maximum accuracy degradation of 5%, still captures almost all of the available speedup and energy efficiency gains.
Figure 7: Approximate Pareto front of the achieved LionHeart mappings, and operational points of both variants of the Harmonica and FLMS methods. For each network, the elbow for LionHeart, i.e., the optimal operating point, is highlighted using a darker hue value.
AlexNet and VGG16, which do not employ separable convolutions, instead present more potential speedup (660% in the full-analog case for VGG16, 550% for AlexNet), due to the higher locality of the employed computations (which leads to less cache misses and memory-induced stalls). Even in those cases, LionHeart can harness a large part of the potential speedup if enough accuracy leeway is allowed. For a 5% accuracy degradation, a 425% speedup was measured in AlexNet and a 285% one in VGG16. AlexNet proved to be the most robust of the investigated networks. Finally, it can be noticed that system-level energy gains closely follow speedups. Indeed, AIMCs tiles are not a major contributor to overall system power (as opposed to processors and memories). Instead, they shorten run-time, ultimately lowering energy requirements. This is achieved both by allowing the execution of MVMs in constant time and by encoding DNN weights in the crossbar structure itself, hence easing the pressure on the memory hierarchy.
Figure 8: For the optimal LionHeart configuration, and both variations of the Harmonica and FLMS methods, (a-e) the MAC ratio (as a percentage) between the number of MACs performed using AIMC and the total number of MACs, and (f-j) the corresponding test set accuracies at $t=$1d. All experiments are repeated $n=5$ times.
TABLE V: Optimized hyper-parameters used for DIANA and Harmonica.
Parameter ResNet8 ResNet20 MobileNet VGG-16 AlexNet
Optimizer SGD + Momentum
Loss Function Cross Entropy
LR Scheduler CosineAnneling
T_max 200
Batch Size 64
Digital
LR 0.0945 0.00184 0.01866 0.0516 0.0157
Momentum 0.86 0.99 0.94 0.88 0.9
Weight Decay 0.00045 0.00017 0.00085 0.00058 0.00011
Analog
LR 0.0617 0.0882 0.0898 0.00619 0.00529
Momentum 0.9 0.82 0.9 0.9 0.99
Weight Decay 0.0008 0.00042 0.00037 0.00037 0.00091
V Comparison to Related Work
We compare against two established heuristic-based methods, DIANA [7] and Harmonica [8]. Comparisons are made using the CIFAR-10 dataset, as this dataset was originally used for evaluation by both of these methods. For DIANA (referred to as First-Last Mapping Strategy (FLMS) to encompass the more generic approach of executing only the first and last layer in high-precision), we consider two scenarios: (i) where the original network is in floating-point precision, akin to LionHeart, and (ii) where the original network is HWA-trained.
For Harmonica, we only consider layer-wise mappings, as channel-wise mappings are too computationally expensive, i.e., they require exploration of a much larger search space dependent on the number of channels rather than the number of layers, to be explored exhaustively. Two scenarios are also considered: (i) where the mapped network is not retrained (Harmonica) and (ii) where the mapped network is retrained using HWA retraining (Harmonica_T). The input to Harmonica is assumed to be HWA-trained. This is problematic, as HWA training requires exhaustive hyper-parameter optimization to achieve optimal results [5] when all layers are implemented using AIMC hardware. This is time demanding and computationally expensive. We avoid comparisons with ODiMO due to their use of a channel-wise split, which markedly reduces the analog utilization, as evidenced by their lower analog mac ratio [22].
To establish strong baselines for downstream comparison to LionHeart, for each DNN architecture, we performed hyper-parameter optimization for HWA training using the Optuna [28] library. Specifically, we employed the Tree-structured Parzen Estimator (TPESampler) and the Median pruning algorithm (MedianPruner) optimization algorithms over 10 trials. The resulting new hyper-parameters are listed in Table V.
FLMS and Harmonica do not determine mappings for different MAC ratio and accuracy values. Consequently, we determine the optimal LionHeart mappings by generating an approximate Pareto front in Fig. 7. Additionally, we determine and compare the average operational point for each method (FLMS and Harmonica). It is observed that the operational points of these methods are captured either on or underneath LionHeat’s Pareto front, meaning that LionHeart discovers the same or even better mapping configurations along its front.
To perform further comparisons using a single operational point in Fig. 8 for each method including LionHeart, we select the elbow for each network. The elbow method transverses operational points along the Pareto front from left-most point (where the quantity encoded using the y-axis is maximized) to the right-most point (where the quantity encoded using the x-axis is maximized). The optimal operation point, or elbow, is distinguished by the fact that before reaching it, the y-axis quantity remains almost unchanged, and after reaching it, rapidly decreases [29].
@@ V Comparison to Related Work
In Fig. 8, it can be observed that for all networks, both variants of FLMS result in a high (near 100%) MAC ratio with zero variance. Although the resulting accuracy values have a small variance, they are unpredictable, i.e., compared to other configurations, the average test accuracy varies greatly across networks. Moreover, for most (4/5) networks, compared to other configurations, the accuracy values are the lowest.
On average, both variants of Harmonica yield higher accuracy values when compared to FLMS, however, the average MAC ratios are low and the variability of both the MAC ratios and accuracy values are larger. When the mapped networks from Harmonica are retrained using HWA retraining, i.e., for Harmonica_T, the average accuracy values increase. It is noted that these networks exhibit similarly variability to the networks originally generated from Harmonica.
LionHeart consistently achieves competitive accuracy values with reduced variation compared to both Harmonica variants, for all networks expect ResNet20, while maintaining relatively high MAC ratios. For the single network (ResNet20) where LionHeart performs sub-optimally, it has a significantly higher MAC ratio with respect to other configurations (not including FLMS). It is speculated that the reason Harmonica does not perform at least as well with respect to accuracy, is that at the beginning, in contrast to LionHeart, it requires an already HWA retrained network, and does not retrain any layers. LionHeart begins with a stronger baseline (a pre-trained floating-point network) and performs retraining, providing an opportunity for accuracy losses to be recovered when each layer is converted to analog.
From Fig. 7, if another point along the Pareto front is taken with a smaller MAC ratio, the resulting test precision significantly improves, that is, for a MAC ratio of 85.63%, the test accuracy is 88.36%. This is larger than the average test accuracy for the other aforementioned configurations.
VI Temporal Drift and Evaluation Time Analysis